OpenStudy (anonymous):

____ 160. How does atomic radius change from top to bottom in a group in the periodic table? a. It first decreases, then increases. b. It tends to decrease. c. It first increases, then decreases. d. It tends to increase. @whpalmer4

OpenStudy (whpalmer4):

have a look at the atomic radius section here: https://en.wikipedia.org/wiki/Periodic_trends

OpenStudy (anonymous):

How many moles of HCl are present in 0.70 L of a 0.33 M HCl solution? a. 0.47 mol b. 0.23 mol c. 0.28 mol d. 0.38 mol

OpenStudy (whpalmer4):

0.33 M means 0.33 moles of solute (HCl) in 1 liter of solution. you've got 0.70 l of solution, so 0.7 l *0.33 mol/1 l = 0.7*0.33 mol =

OpenStudy (anonymous):

0.231

OpenStudy (whpalmer4):

0.231 mol

OpenStudy (anonymous):

yeah